OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. : x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

OpenStudy (anonymous):

@RowanMartinFan

OpenStudy (anonymous):

@Makoto_Tachibana

OpenStudy (anonymous):

mkay i was tagged

OpenStudy (anonymous):

960, do i have to explain?

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

if it was right give me a medal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are u sure ur right or u guesing

OpenStudy (anonymous):

x is the total of number people who downloaded the music 1x = 0.6x + 384 1x - 0.6x = 384 0.4x = 384 0.4x/0.4 = 384/0.4 x = 960 Now you have to look back at the question, which says that 60% of people downloaded this song. Simply take 60% of the value of x. 960 x 0.6 = 576.

OpenStudy (anonymous):

so its 576 not 960??

OpenStudy (anonymous):

oops yes sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill post another q

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

but gtg